The breadth of the room is one third of its length. If the parimeter of the room is 64 daimeter, find length and breath of the room.
Answer by nerdybill(7384)

The breadth of the room is one third of its length. If the parimeter of the room is 64 daimeter, find length and breath of the room.
.
Think of "breath" as "width".
.
Let x = length
then
x/3 = width
.
2(x + x/3) = 64
x + x/3 = 32
3x + x = 96
4x = 96
x = 24 (length)
.
breath:
x/3 = 24/3 = 8 (breath)